Hi there.
The problem I need help with is to solve algebraically for "n" in
nP4=8(n-1P3).
Thanks

Answer by Theo(13342)

general way to solve this problem is as follows:
np4 = 8(n-1p3)
simplify by removing parentheses to get:
np4 = 8n - 8p3
subtract 8n from both sides of the equation to get:
np4 - 8n = - 8p3
factor out the n to get:
n(p4-8) = -8p3
divide both sides of the equation by (p4-8) to get:
n = -8p3 / (p4-8)